Find the slope of a line that is perpendicular to the line with the equation 5x+6y=14
Answer by mananth(16946)

5 x + 6 y = 14
Find the slope of this line

6 y = -5 x + 14
Divide by 1
y = - 5/ 6 x + 7/3
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 5/6

The slope of a line perpendicular to the above line will be the negative reciprocal 6/5
m1*m2=-1